the probability that a computer crashes during a week is 0.18. Room 12 has 6 computer in it. what is the probability that none of the computer crashes during the first week?
Found 2 solutions by flame8855, lynnlo:
Answer by flame8855(424)

probasbility of not crashing = 1-0.18=0.82
0.82^3= 0.55

Answer by lynnlo(4176)

-a^3b^2c^2-a^2b^3c^2+a^b^2c^3